# Experiences with Sub-Arctic Sensor Network Deployment

### v2

@@ V Results and Discussions
We were well aware that energy harvesting could have been a potential solution to the battery lifetime issue. However, we were hesitant to depend on it for several reasons: first, there were no off-the-shelf, proven, and reliable energy harvesting solutions available on the market. Countless energy harvesting architectures exist in the literature, but most are experimental. Though we could have experimented with energy harvesters, we deliberately avoided this considering the immediate requirement for data. The availability of high energy density primary cells in small form factor made the choice easy. Another factor in favour of the battery is the lack of reliable ambient energy sources. While solar energy is the preferred option for outdoor deployments, the harsh Icelandic winter weather renders it largely inefficient. Nonetheless, over the course of the project, we have uncovered a unique source of energy for outdoor deployments, which is the temperature difference between the soil and the air. While this energy is particularly available at ForHot due to the geothermal activity, it is present even in places without geothermal heating [7]. One notable feature that sets this energy source apart from solar is its near 24/7 availability, providing power both day and night. However, its power density is often limited, making a multisource energy harvester that combines both solar and soil-air thermal energy the optimal choice for outdoor deployment in harsh conditions. We have advanced the research in this context and have developed prototypes and proof-of-concept devices that work with soil-air thermal energy [8, 7] and multisource energy harvesting architectures [9].
COVID-19 had a substantial impact on development and deployment. In addition to travel restrictions, the inability to access the lab environment significantly delayed the work. Moreover, the global chip shortage nearly jeopardized our hardware architecture due to the unavailability of CMWX1ZZABZ-071 modules. However, we managed to refurbish and reuse OCTA boards from a different project, highlighting the importance of reusable hardware. Additionally, during the lockdowns, we were able to carry out limited small-scale deployments in a controlled outdoor environment. This helped us to uncover some of the inefficiencies of the initial hardware design. For instance, no external watchdog timer was initially included in the design as we relied on the internal one. However, during the test run, a couple of sensor nodes abruptly ceased to function and required a hard restart by removing the battery. Upon inspection, it was discovered that the firmware had entered a faulty state due to bugs in the DASH7 stack. Therefore, an external watchdog was added to later versions to handle any unexpected firmware failures.
@@ VI Conclusion
We have discussed the design, deployment and maintenance of low-power wireless sensor networks for monitoring and sensing at the ForHot sites in Iceland. The network has attained more than three years of lifetime and is still functional. While the primary cells provide a quick solution, it requires further scrutiny to evaluate their long-term performance.

### v3

@@ V Results and Discussions
Feasibility of Energy Harvesting: We were well aware that energy harvesting could have been a potential solution to the battery lifetime issue. However, we were hesitant to depend on it for several reasons: first, there were no off-the-shelf, proven, and reliable energy harvesting solutions available on the market. Though we could have experimented with energy harvesters, we deliberately avoided this, considering the immediate requirement for data. The availability of high energy density primary cells in small form factor made the choice easy. Another factor favoring the battery is the lack of reliable ambient energy sources. Solar energy is the preferred option for outdoor deployments; however, the harsh Icelandic winter weather makes it largely inefficient. Nonetheless, over the course of the project, we have uncovered a unique source of energy for outdoor deployments, which is the temperature difference between the soil and the air. While this energy is particularly available at ForHot due to the geothermal activity, it is present even in places without geothermal heating [10]. One notable feature that sets this energy source apart from solar is its near 24/7 availability, providing power both day and night. However, its power density is often limited, making a multi-source energy harvester that combines both solar and soil-air thermal energy the optimal choice for outdoor deployment in harsh conditions. We have advanced the research in this context and have developed prototypes and proof-of-concept devices that work with soil-air thermal energy [10] and multi-source energy harvesting architectures [11].
@@ VI Conclusion
We have discussed the design, deployment and maintenance of low-power wireless sensor networks for environment monitoring and sensing at three remote sites in Iceland. The networks, functional since 2021, have achieved a lifespan of over three years with sensor nodes powered by batteries and remain functional. The mid-range communication protocol D7A is proven to be a reliable choice, particularly due to its low-power and efficient downlink capabilities.
Acknowledgement
The authors thank Dr. Páll Sigursson, Prof. Bjarni Dirik Sigursson, other FutureArctic colleagues, and the staff at the Agricultural University of Iceland, Hverageri for their enormous support and assistance during the deployment and maintenance of the network. The authors also thank the anonymous reviewers of IEEE WCNC’25 and EWSN’24 for their constructive feedback and comments on our work.
This research has received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Sokolowski-Curie grant agreement No 813114.